An object of mass 1kg moving on a horizontal surface with initial velocity 8m/s comes to rest after 10s . If one wants to keep the object moving on the same surface with velocity 8m/s the force required is....1. 0.4N2. 0.8N3. 1.2N4. Zero

Grade:11

2 Answers

Diwankar singh
6 years ago
mass=1kginitial velocity`u`=8m/sfinal velocity`v`=0m/stime=10 sec.therefore, acceleration=v-u/t =0-8/10 =-0.8m/s^2Force=mass*acceleration =1*(-0.8) = -0.8NWe know that force cannot be in negative therefore, force applied =0.8N
nidhi
6 years ago
v = u + at
v = 0
0 = 8 – \displaystyle \mug(10)
\mu = 8/100 = 0.08
To move the body withconstant velocity on this surface, internal force applied should be equal to friction force
F = \mumg
   = (0.08)(1)(10)
   = 0.8N